OpenStudy (anonymous):

Given a cone with a volume of 288 inches cubed and height 7 in., find the base radius of the cone. Use 3.14 for pi. Round your answer to the tenths place. a. 2.4 in c. 3.5 in b. 2.8 in d. 6.3 in

OpenStudy (anonymous):

Volume of a cone is (1/3)*pi*(r^2)*h 288 = (1/3) * 3.14 * (r^2) * 7 r^2 = 39.3 square root both sides r=6.3 d)6.3 in

OpenStudy (anonymous):

thank you !!!!

OpenStudy (jhannybean):

For future reference, I would just leave \(\pi\) as \(\pi\) and no restrict it to 3.14

OpenStudy (anonymous):

you're welcome :)

OpenStudy (anonymous):

okay thanks for the tip

OpenStudy (jhannybean):

No problem :)